OpenStudy (anonymous):

write an equation in slope-intercept form for the line that passes through the given point and is perpendicular to the graph of the equation. (4,-2),y=3x+5

OpenStudy (anonymous):

first off, where are you getting stuck at? ^_^

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

no problem first lets figure out what the slope would be for a line perpendicular to the line we have been given, so find 'm' out of your line equation of y=3x+5, then multiply it by (-1) then flip that fraction so for example if m=8, then I would do (-1)*8= -8, then flip it: -1/8 ^_^ can you do this procedure for your m?

OpenStudy (anonymous):

y=-3/1-5

OpenStudy (anonymous):

you got the slope for the perpendicular line right ^_^

OpenStudy (anonymous):

next step is to plug in new m into the slope-intercept form y = mx+b y=(-1/3)x+b now we need to plug in the point the problem gave us, and solve for b

OpenStudy (anonymous):

sorry, m = -1/3 not -3/1

OpenStudy (texaschic101):

y = 3x + 5 perpendicular line will be negative reciprocal of the slope. So, if the slope is 3, the negative reciprocal of the slope is -1/3. What you do is " flip " the slope and change the sign.

OpenStudy (anonymous):

@texaschic101 could i get you to look at another problem i'm trying to work out?

OpenStudy (texaschic101):

sure....I can't make no promises that I can do it, but I will look at it

OpenStudy (anonymous):

so would it be 3/1

OpenStudy (texaschic101):

no....we know that the slope we need to use is -1/3. We will now use this formula : y - y1 = m(x - x1) slope(m) = -1/3 (4,-2) x1 = 4 and y1 = -2 now lets sub y - (-2) = -1/3(x - 4) -- distribute through the parenthesis y + 2 = -1/3x + 4/3 -- subtract 2 from each side y = -1/3x + 4/3 - 2 y = -1/3x + 4/3 - 6/3 y = -1/3x - 2/3 ==> your perpendicular line any questions ?

OpenStudy (anonymous):

then how do u do (-5,2) y=1/2x-3 because i did that one so different than you did

OpenStudy (texaschic101):

you looking for parallel or perpendicular ?

OpenStudy (anonymous):

perpendicular

OpenStudy (texaschic101):

the slope of the line given is 1/2. But we need perpendicular slope...so flip the slope and change the sign making the slope we need -2. you can either do it using this formula : y - y1 = m(x - x1) or you can just use y = mx + b (1) y = mx + b (-5,2) x = -5 and y = 2 slope(m) = -2 now sub and find b 2 = -2(-5) + b 2 = 10 + b 2 - 10 = b -8 = b so your equation is : y = -2x - 8 (2) y - y1 = m(x - x1) slope(m) = -2 (-5,2) x1 = -5 and y1 = 2 now lets sub y - 2 = -2(x -(-5) y - 2 = -2(x + 5) y - 2 = -2x - 10 y = -2x - 10 + 2 y = -2x - 8 you can do it either way

OpenStudy (anonymous):

okay thanks